OpenStudy (ttop0816):

please help! cos(pi/3+b) 1/2(cosb-root3sinb) 1/2(cosb+root3sinb) root3(cosb-1/2sinb) root3(cosb+1/2sinb)

Parth (parthkohli):

This?\[\cos\left(\dfrac{\pi}{3} + b\right)\]

Parth (parthkohli):

If so, in that case, use\[\cos(a + b) = \cos (a)\cos(b) - \sin(a)\sin(b)\]

OpenStudy (ttop0816):

yes i got it till that part!

Parth (parthkohli):

Also remember that \(\cos\left(\pi/3\right) = 1/2\) and \(\sin(\pi/3) = \sqrt3 /2\). That is all you need!

OpenStudy (ttop0816):

yes i remember! but the problem that i have is the last part of solving ): the calculation part; im kinda confused!

Parth (parthkohli):

OK. Let's do this!

Parth (parthkohli):

\[\cos\left(\dfrac{\pi}{3} + b\right) = \cos \left(\dfrac{\pi}{3} \right)\cos(b) - \sin\left(\dfrac{\pi}{3}\right)\sin(b)\]

Parth (parthkohli):

It really can't be hard from here. Can it?

OpenStudy (ttop0816):

then it would be (1/2) cos(b) - (root3) sin (b)??

Parth (parthkohli):

Yup, except that is \(\dfrac{\sqrt3}{2}\)

OpenStudy (ttop0816):

@ParthKohli thank you so much!!

Parth (parthkohli):

You're welcome!